OpenStudy (anonymous):

Quick help please! Using kite LMON, solve for x. 4 7 21 cannot be determined

OpenStudy (anonymous):

@warningiflirt please another one

OpenStudy (anonymous):

6x+17=9x-4 6x-9x=-4-17 -3x=-21 x=7

OpenStudy (anonymous):

HAHA your awesome i got the same thing @warningiflirt

OpenStudy (anonymous):

thanks :)